Convert into rectangular coordinates (4,-2π\3)

1
Expert's answer
2022-06-06T13:51:06-0400
"x=r\cos \theta=4\cos(-2\pi/3)=-2"

"y=r\sin \theta=4\sin(-2\pi/3)=-2\sqrt{3}"



"(-2, -2\sqrt{3})"